OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1). y + 3 = −5(x − 2) y − 2 = −5(x + 3) y + 3 = −one fifth(x − 2) y − 2 = −one fifth(x + 3)

OpenStudy (anonymous):

@phi last one

OpenStudy (anonymous):

Yes, that is correct!

OpenStudy (anonymous):

u sure your not phi

OpenStudy (anonymous):

I'm definitely sure of 2 things: the correct answer, and the fact that I'm not phi!

OpenStudy (anonymous):

lol ok

OpenStudy (phi):

the first step is find the slope. can you do that ?

OpenStudy (anonymous):

Good luck in all of your studies and thanks for the recognition! @ryanmj6

OpenStudy (anonymous):

idk how to find the slope sorry

OpenStudy (anonymous):

brb my girlfriend needs help with something

OpenStudy (anonymous):

By "last one", did you mean this is your last question, or that you are selecting the last choice for your answer? I thought you were picking your answer.

OpenStudy (anonymous):

last answer

OpenStudy (anonymous):

i guess but it was supposed to be last question

OpenStudy (texaschic101):

find the slope first slope (m) = (y2 - y1) / (x2 - x1) (-3,2) -- x1 = -3, y1 = 2 (2,1) -- x2 = 2, y2 = 1)

OpenStudy (anonymous):

The slope, which is referred to as "m" is as follows:\[m = \frac{ y_{2} - y _{1} }{ x _{2} - x _{1} }\]where\[(x _{1}, y _{1})\]and\[(x _{2}, y _{2})\]are your given points. So, you would have:\[m = \frac{ 2 - 1 }{ -3 - 2 }\]